OpenStudy (anonymous):

prove the identity cos(3x)=cos^3(x)-3sin^2(x)cos(x) I think I understand a lot of it...but I need help though.

OpenStudy (anonymous):

try using cos (a + b ) where a = x and b = 2x

OpenStudy (anonymous):

so i should focus on the LHS? @chandanjha

OpenStudy (anonymous):

just use the identity cos(a+b) = sina sinb+cosa cob

OpenStudy (anonymous):

there should be a minus sign cross check it

OpenStudy (campbell_st):

start with the left and write it as \[\cos(2x + x)\] then use the expansion \[\cos(2x)\cos(x) - \sin(2x)\sin(x)\] next use the double angle information such as sin(2x) = 2sin(x)cos(x) and then its just processing to the answer on the right

OpenStudy (anonymous):

in between sinb and cosa @chandanjha

OpenStudy (anonymous):

use as @campbell_st has suggested

OpenStudy (anonymous):

its cos a cos b - sin a sin b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next use the double angle information such as sin(2x) = 2sin(x)cos(x) i dont get this

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

i wrote out the expansion but what do i dowith it?

OpenStudy (campbell_st):

this is a standard expansion of sin(2x) sin(2x) = 2sin(x)cos(x) so just substitute it and you get cos(2x)cos(x) -2sin(x)cos(x)sin(x) which can be written as \[\cos(2x)\cos(x) - 2\sin^2(x)\cos(x)\] next you need to look at the expansion of cos(2x) there are several versions but a good starting point is \[\cos(2x) = \cos^2(x) - \sin^2(x)\] so now substitute this into \[\cos(2x)\cos(x) - 2\sin^2(x)\cos(x)\]

OpenStudy (anonymous):

when you expand sin@x i dont see where the heck cos comes from

OpenStudy (anonymous):

i dont see any identity...

OpenStudy (campbell_st):

its this sin(2x) can be written as sin(x + x) using the expansion for the sum of 2 angles in sin, you get sin(x +x) = sin(x)cos(x) + sin(x)cos(x) = 2sin(x)cos(x) if you look at cos cos(2x) = cos(x +x) = cos(x)cos(x) - sin(x)sin(x) = cos^2(x) - sin^2(x)

OpenStudy (anonymous):

ooooh so you have to like backtrack to go forward again?

OpenStudy (campbell_st):

not really when studying this topic of trig equations, you start with sum and difference and then go to the double angle expansions.... some people memorise them, I just use the sum and difference but have the same angles... its just stuff you need to know... to succeed in trig equations.

OpenStudy (anonymous):

so I substituted it in, and got\[\cos ^{2}x-\sin^2xcosx-2\sin^2xcosx\] right?

OpenStudy (primeralph):

Looks like you're in good hands. Good luck.

OpenStudy (campbell_st):

not quite right \[(\cos^2(x) - \sin^2(x))\cos(x) - 2\sin^2(x)\cos(x)\] which becomes \[\cos^3(x) - \sin^2(x)\cos(x) - 2\sin^2(x)\cos(x)\] collect the like terms and you'll get the right hand side...

OpenStudy (anonymous):

why is cos to the third?

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

nvm

OpenStudy (campbell_st):

well you have to distribute cos(x) \[(\cos^2(x) - \sin^2(x))\cos(x) = \cos^2(x) \times \cos(x) - \sin^2(x)\times \cos(x)\] which gives \[\cos^3(x) - \sin^2(x)\cos(x)\]

OpenStudy (anonymous):

thank youuuu!

OpenStudy (campbell_st):

These questions are just tedious.... just remember the sum and difference expansions and you can then work out the double angle information. good luck

OpenStudy (anonymous):

thanks